whats the equation of the line that passes through point (-2, -3) and has a slope of 5.
Answer by lwsshak3(11628)

whats the equation of the line that passes through point (-2, -3) and has a slope of 5.
**
Standard form of an equation for a straight line: y=mx+b, m=slope, b=y-intercept
For given equation:
y=5x+b
solving for b using given point ,(-2, -3),thru which given line passes
-3=5*-2+b
-3=-10+b
b=7
equation: y=5x+7x